OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 6x – y = 1 4x – 3y = –11 4(–6x + 1) – 3y = –11 4(6x + 1) – 3y = –11 4x – 3(6x – 1) = –11 4x – 3(–6x – 1) = –11

OpenStudy (anonymous):

C

OpenStudy (anonymous):

6x-y=1 6x-1=y(Jusy rearrange terms in this way) Substitute this in the second equation... 4x-3(6x-1) = -11